OpenStudy (anonymous):

Write the equation for the line tangent to the circle x^2 + y^2 = 29 at the point (2, 5).

OpenStudy (anonymous):

do you have your derivative?

OpenStudy (anonymous):

this is all i'm given

OpenStudy (anonymous):

do you know how to calculate the derivative implicitly?

OpenStudy (anonymous):

no..

OpenStudy (jit4won):

find equation of radius

OpenStudy (anonymous):

i'm sorry... is this a calculus problem?

OpenStudy (jit4won):

then use it's slope to find equation of tangent

OpenStudy (anonymous):

its algebra 2b. I don't believe its calculus

OpenStudy (anonymous):

then @jit4won , has the right approach... sorry..

OpenStudy (jit4won):

thanks

OpenStudy (jit4won):

equation of radius = 5x-2y= 0

OpenStudy (jit4won):

equation of tangent = 2x+5y =29

OpenStudy (jit4won):

Anything else that I may help you with ?

OpenStudy (anonymous):

Actually, yeah!! Write an equation in standard form for the hyperbola with vertex (4, 0), and focus (10, 0).

OpenStudy (jit4won):

focus =(a.0) =10,0 a= 10

OpenStudy (jit4won):

sorry vertex = a,0

OpenStudy (anonymous):

I'm told the answer is supposed to be in the form of a fractional equation

OpenStudy (jit4won):

yes

OpenStudy (anonymous):

I'm lost

OpenStudy (jit4won):

1 min

OpenStudy (jit4won):

vertex = 4,0 a= 4 focus = 10.0 c= 10 c^2 = a^2 +b^2 b^2= 10^2-4^2= 84 equation = x^/16 -y^2/84 = 1

OpenStudy (anonymous):

awesome!

OpenStudy (jit4won):

Thank you very much

OpenStudy (anonymous):

thank you!